# Milnor number of an invariant singularity: generalization of Chulkov’s inequality

### v3

@@ Abstract.
S.P. Chulkov has proven that the Milnor number of a function germ $f$ with zero 2-jet invariant with respect to a nontrivial action of $\mathbb{Z}_{p}$ is at least p1p-1 for prime $p$. In this paper we prove a generalization of this inequality for actions of arbitrary abelian groups using M. Roberts’ results on invariant morsifications for invariants of real group actions.
@@ 2. Definitions and auxiliary statements
In a suitable local system of coordinates any finite group action on $(\mathbb{C}^{n},0)$ can be linearized (see [6]). So we wil consider all group actions to be induced by a linear representation of $G$. Furthermore, any linear representation of a finite abelian group is diagonalazable, hence further we will consider only actions of the form $g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ with $\chi_{i}$ being group characters of $f$.
@@
For a function germ $f$ $J_{f}$ will denote the Jacobian ideal $\frac{\partial f}{\partial x_{1}},\dots,\frac{\partial f}{\partial x_{n}}>$. $Q_{f}$ is the Jacobian algebra $\mathbb{C}{x_{1},\dots,x_{n}}/J_{f}$. $Q_{f}$ always admits a monomial basis, i.e. the basis consisting of monomials
@@
A group action on $(\mathbb{C}^{n},0)$ on is called real if it admits an invariant quadratic form of rank $n$.
@@
Consider the action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow\\ \longrightarrow(x_{1},\dots,x_{k},\chi_{k+1}(g)x_{k+1},\dots,\chi_{n}(g)x_{n}),\chi_{k+1},\dots,\chi_{n}\neq 1$ and an invariant function $f$ with a zero 2jet2-jet and an isolated critical point at the origin.
@@ 3. Proof of main theorem
For a group action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ we will denote $\tau_{\mathbb{R}}$ the action on $\mathbb{C}^{2n}$ defined as $g\circ(x_{1},\dots,x_{n},y_{1},\dots,y_{n})\longrightarrow\\ (\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n},\bar{\chi}_{1}(g)y_{1},\dots,\bar{\chi}_{n}(g)y_{n})$.
@@
For a function $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0) ff$$\oplus f$ is the function $f(x_{1},\dots,x_{n})+f(y_{1},\dots,y_{n})$.
@@
If $p_{1},\dots,p_{\mu}$ form a monomial basis for the Jacobian algebra of $f$ then $p_{i}(x)p_{j}(y),1\leq i,j\leq\mu$ form a monomial basis in the Jacobian algebra of $f\oplus f$.
@@
Start with an arbitrary monomial basis $q_{1}(x,y),\dots,q_{\mu^{2}}(x,y)$} for $Q_{f\oplus f}$. Each $q_{i}$ factors as as a product of monomials in $x$ and $y$: $q_{i}=q_{i1}(x)q_{i2}(y)$. For each $i$ $q_{i1}\notin J_{f}$, as otherwise $q_{i}\in J_{f\oplus f}$. Hence we have a decomposition $q_{i1}=\Sigma_{j}c_{ij}p_{j}(x)$. Similarly, each $q_{i2}$ is a linear combination of $p_{j}(y)$. Therefore $p_{i}(x)p_{j}(y)$ form a system of generators for $Q_{f\oplus f}$. As there are exactly $\mu^{2}$ of them, they form a basis.
@@
$\nu(f\oplus f)\geq\mu(f)$.
@@
Consider a monomial basis $p_{i}(x)p_{j}(y),1\leq i,j\leq\mu$ of $Q_{f\oplus f}$ as described in the previous theorem. Each $p_{i}(x)$ is an equivariant function with respect to the action $\tau$ with some weight $\lambda$. As $G$ acts by conjugate characters on $y$, $p_{i}(y)$ is equivariant with the inverse weight $\bar{\lambda}$. Hence  $\forall\;1\leq i\leq\mu\;p_{i}(x)p_{i}(y)\in Q^{G}_{f\oplus f}$.
Now we can conclude the proof of the main theorem. As $\tau_{\mathbb{R}}$ is a real representation, we can use Roberts’ inequality for the Milnor number of $f\oplus f$:
@@ 4. Tightness of the bound
The next step is to produce examples showing the tightness of this bound. The trivial example is the 1-dimensional representation: if a character $\chi$ of group $G$ has order $k$ and $G$ acts on $\mathbb{C}$ via this character, the invariant function $x^{k}$ has the lowest possible Milnor number k1k-1. However, this is unsatisfactory, as we would ideally like to prove that a)the bound is tight in arbitrarily large dimensions; b)the Milnor number is bounded by the length of the shortest orbit when there are orbits of different lengths. To show this, we present the following example.
Let $n$ be an odd positive integer, $g$ a generator of a group $\mathbb{Z}_{2^{n}+1}$, $h$ a generator of $\mathbb{Z}_{3}$, $\epsilon,\sigma$ – primitive roots of unity of order $2^{n}+1,3$, respectively .
$\mathbb{Z}_{2^{n}+1}\times\mathbb{Z}_{3}$ acts on $\mathbb{C}^{n}$ in the following manner: $g\circ(x_{1},\dots,x_{n})$
$(\epsilon^{(-2)}x_{1},\epsilon^{(4)}x_{2},\dots,\epsilon^{(-2)^{n-1}}x_{n-1},\epsilon x_{n})$, $h\circ(x_{1},\dots,x_{n})=\sigma(x_{1},\dots,x_{n})$.
The polynomial $x_{1}^{2}x_{2}+x_{2}^{2}x_{3}+\dots+x_{n-1}^{2}x_{n}+x_{n}^{2}x_{1}$ is invariant and has Milnor number $2^{n}$, the length of the shortest orbit (e.g. of the point $(0,0,\dots,0,1)$) is $2^{n}+1$, however the longest orbit (e.g. of $(1,1,\dots,1)$) has length $3(2^{n}+1)$

### v4

@@ Abstract.
We prove a lower bound for the Milnor number of function germ invariant with respect to a finite abelian group action. It is shown that this bound is tight for functions of arbitrarily many variables. We also prove the function germs that reach this lower bound are equivariantly stable, i.e. invariant analogues of Morse singularities.
@@ 2. Definitions and auxiliary statements
In a suitable local system of coordinates any finite group action on $(\mathbb{C}^{n},0)$ can be linearized (see [6]). So we wil consider all group actions to be induced by a linear representation of $G$. Furthermore, any linear representation of a finite abelian group is diagonalazable, hence further we will consider only actions of the form $g\circ(x_{1},\ldots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\ldots,\chi_{n}(g)x_{n})$ with $\chi_{i}$ being group characters of $f$.
@@
For a function germ $f$ $J_{f}$ will denote the Jacobian ideal $\frac{\partial f}{\partial x_{1}},\ldots,\frac{\partial f}{\partial x_{n}}>$. $Q_{f}$ is the Jacobian algebra $\mathbb{C}\{x_{1},\ldots,x_{n}\}/J_{f}$. $Q_{f}$ always admits a monomial basis, i.e. the basis consisting of monomials
@@
A group action on $(\mathbb{C}^{n},0)$ is called real if it admits an invariant quadratic form of rank $n$.
@@
Consider the action $\tau:g\circ(x_{1},\ldots,x_{n})\longrightarrow\\ \longrightarrow(x_{1},\ldots,x_{k},\chi_{k+1}(g)x_{k+1},\ldots,\chi_{n}(g)x_{n}),\chi_{k+1},\ldots,\chi_{n}\neq 1$ and an invariant function $f$ with a zero 2jet2-jet and an isolated critical point at the origin.
@@ 3. Proof of main theorem
For a group action $\tau:g\circ(x_{1},\ldots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\ldots,\chi_{n}(g)x_{n})$ we will denote $\tau_{\mathbb{R}}$ the action on $\mathbb{C}^{2n}$ defined as $g\circ(x_{1},\ldots,x_{n},y_{1},\ldots,y_{n})\longrightarrow\\ (\chi_{1}(g)x_{1},\ldots,\chi_{n}(g)x_{n},\bar{\chi}_{1}(g)y_{1},\ldots,\bar{\chi}_{n}(g)y_{n})$.
@@
For a function $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0) ff$$\oplus f$ is the function $f(x_{1},\ldots,x_{n})+f(y_{1},\ldots,y_{n})$.
@@
If $p_{1},\ldots,p_{\mu}$ form a monomial basis for the Jacobian algebra of $f$ then $p_{i}(x)p_{j}(y),1\leq i,j\leq\mu$ form a monomial basis in the Jacobian algebra of $f\oplus f$.
@@
Start with an arbitrary monomial basis $q_{1}(x,y),\ldots,q_{\mu^{2}}(x,y)$} for $Q_{f\oplus f}$. Each $q_{i}$ factors as as a product of monomials in $x$ and $y$: $q_{i}=q_{i1}(x)q_{i2}(y)$. For each $i$ $q_{i1}\notin J_{f}$, as otherwise $q_{i}\in J_{f\oplus f}$. Hence we have a decomposition $q_{i1}=\Sigma_{j}c_{ij}p_{j}(x)$. Similarly, each $q_{i2}$ is a linear combination of $p_{j}(y)$. Therefore $p_{i}(x)p_{j}(y)$ form a system of generators for $Q_{f\oplus f}$. As there are exactly $\mu^{2}$ of them, they form a basis.
@@
$\nu(f\oplus f)=k_{1}^{2}+\ldots+k_{m}^{2}$, with $k_{i}$ being the multiplicities of different one-dimensional representations in the irreducible decomposition of $Q_{f}$.
@@
As proven in the previous theorem $Q_{f\oplus f}$ is isomorphic ( as a vector space) to the tensor product of $Q_{f}$ with itself. On the second factor of this tensor product (i.e. on monomials in $y_{i}$) $G$ acts by conjugate characters, so the action on the group on the second factor can be identified with the action on the conjugate space $Q^{*}_{f}$. According to a well-known result in representation theory (see, e.g., [7], pp. 433–434) for each finite-dimensional representation $V$ the representation $V\otimes V^{*}$is naturally isomorphic to $Hom(V,V)$, and the invariant subspace of $V\otimes V^{*}$ is isomorphic to the space of equivariant endomorphisms of $V$. The dimension of the space of equivariant endomorphisms is $k^{2}_{1}+\ldots+k^{2}_{m}$ with $k_{i}$ being the multiplicities of irreducible representations in the decomposition of $V$.
Since $\mu(f)=k_{1}+\ldots+k_{m}$, we have, in particular, $\nu(f\oplus f)\geq\mu(f)$. Now we can conclude the proof of the main theorem. As $\tau_{\mathbb{R}}$ is a real representation, we can use Roberts’ inequality for the Milnor number of $f\oplus f$:
@@ 4. Tightness of the bound
The next step is to produce examples showing the tightness of this bound. The trivial example is the 1-dimensional representation: if a character $\chi$ of group $G$ has order $k$ and $G$ acts on $\mathbb{C}$ via this character, the invariant function $x^{k}$ has the lowest possible Milnor number k1k-1. However, this is unsatisfactory, as we would ideally like to prove that the bound is tight in arbitrarily large dimensions. To show this, we present the following examples.
Let $n$ be an odd positive integer, $d_{1},\ldots,d_{n}\geq 2$ – natural numbers, $g$ – generator of a group $\mathbb{Z}_{d_{1}\ldots d_{n}+1}$, $\epsilon$ – pimitive root of unity $1$ of degree $d_{1}\ldots d_{n}+1$. $\mathbb{Z}_{d_{1}\ldots d_{n}+1}$ acts on $\mathbb{C}^{n}$ in the following fashion:
$g\circ(x_{1},\ldots,x_{n})=(\epsilon x_{1},\epsilon^{-d_{1}}x_{2},\ldots,\epsilon^{(-1)^{n-2}d_{1}\ldots d_{n-2}}x_{n-1},\epsilon^{(-1)^{n-1}d_{1}\ldots d_{n-1}}x_{n})$
The polynomial $x_{1}^{d_{1}}x_{2}+x_{2}^{d_{2}}x_{3}+\ldots+x_{n-1}^{d_{n_{1}}}x_{n}+x_{n}^{d_{n}}x_{1}$ is invariant with respect to this action. It has an isolated critical point, as it is an invertible polynomial of the loop type (see., e.g., [8, 9]), and its Milnor number is equal to $d_{1}\ldots d_{n}$.
The length of any nontrivial orbit in this case is $d_{1}\ldots d_{n}+1$, so the bound $\mu\geq l(\tau)-1$ is tight.
5. Equivariantly stable germs
There are many papers on the classification of equivariantly simple singularities (see., e.g., [10], [11], [12]). Roughly speaking, an equivariantly stable singularity is the simplest of the equivariantly simple singularities, i.e. the one that is only adjacent to itself. If the case of real group action the equivariantly stable singularity is the nondegenerate one, as the deformation of a Morse function is again a Morse function. For the formal definition of equivariant stability see below.
We will denote $\mathfrak{m}_{G}$ the maximal ideal in the ring of $G$-invariant analytic germs. Since our action has no fixed points outside of origin, every element of $\mathfrak{m}_{G}$ has a critical point at the origin. Consider $\mathcal{O}^{r}_{G}$ — the space of $r$-jets of elements in $\mathfrak{m}_{G}$. The group $D^{r}_{G}$ of $r$-jets of $G$-equivariant diffeomorphisms $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{n},0)$ acts on this space.
A germ $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ is equivariantly stable if for all large enough $r$ $D^{r}_{G}$-orbit of $r$-jet of $f$ in $\mathcal{O}^{r}_{G}$ is open.
Theorem 5.1.
$f$ is equivariantly stable iff $\nu(f)=1$.
Proof.
Let $f$ be equivariantly stable. As the orbit of $r$-jet of $f$ is open for large enough $r$ the tangent space to the orbit has dimension equal to the dimension of $\mathcal{O}^{r}_{G}$ itself. The tangent space to the orbit at $r$-jet of $f$ is the projection of the jacobian ideal of $f$ to $\mathcal{O}^{r}_{G}$. But if for all $r$ large enough the projection of $J_{f}$ covers all $\mathcal{O}^{r}_{G}$ $J_{f}$ contains $\mathfrak{m}_{G}$. Since $f$ has a critical point at the origin $J_{f}$ does not contain 1 and $\nu(f)=1$.
Now assume $\nu(f)=1$. The Milnor number of $f$ is finite, as the invariants of a finite group action have the separating property, i.e. for any two orbits of the $G$-action there is an invariant function that is equal to 1 on the first orbit and 0 on the second. But if $\nu(f)=1$, then $J_{f}=\mathfrak{m}_{G}$ and any invariant function takes the same value at any point of the singular locus of $f$ as it does at the origin. So the separating property fails if the critical point of $f$ is non-isolated. Since $f$ has a finite Milnor number we can apply the Slodowy theorem (see [13]), according to which the invariant versal deformation of $f$ is equal to $f+\lambda_{1}h_{1}+\ldots+\lambda_{\nu}h_{\nu}$, $h_{j}$ — linearly independent invariant elements of $Q_{f}$. Since $\nu(f)=1$, the versal deformation of $f$ is trivial, i.e. has the form $f_{\lambda}=f+\lambda$, and the equivariant stability follows immediately. ∎
In particular for the nondegenerate singularity $\mu(f)=\nu(f)=1$, as $Q_{f}$ is equal to $\mathbb{C}$.
Theorem 5.2.
If $\mu(f)=l(\tau)-1$ $f$ is equivariantly stable
Proof.
If $\mu(f)=l(\tau)-1$, the inequality $\mu^{2}(f)\geq(\mu(f)-1)l(\tau)+1$ becomes an equality. Since also $\mu^{2}(f)\geq(\nu(f\oplus f)-1)l(\tau)+1$ and $\nu(f\oplus f)\geq\mu(f)$, it can only become an equality if $\nu(f\oplus f)=\mu(f)$. As was proven earlier, $\nu(f\oplus f)=k_{1}^{2}+\ldots+k_{m}^{2}$, with $k_{i}$ being the multiplicities of one-dimensional representations in $Q_{f}$. With $\mu(f)$ equal to $k_{1}+\ldots+k_{m}$, the equality $\nu(f\oplus f)=\mu(f)$ can only hold if $\forall\;i\;k_{i}=1$. In particular, the multiplicity of the trivial representation (equal to $\nu(f)$) is 1. ∎